### v1

# A study of gamma-ray emission from OJ 287 using Fermi-LAT from 2015-2023

Abstract
We search for gamma-ray emission from OJ287 in the energy range from 0.1-300 GeV during 2015-2023, in coincidence with an extensive observing campaign to monitor the optical flux variability and polarization, as discussed in arXiv:2311.02372. We present results for eight segments in the aforementioned period, with each segment corresponding to an observing season. We find non-zero gamma-ray emission for all the eight segments. The photon and energy flux observed during this period is $\sim 5.2\times 10^{-8}\rm{ph~{}cm^{-2}~{}s^{-1}}$ and $2.75\times 10^{-5}\rm{MeV~{}cm^{-2}~{}s^{-1}}$, respectively, while the SED can be fitted with a power law with a slope of $2.12\pm 0.02$. The observed luminosity is between $10^{45}-10^{46}$ ergs/sec.
1 Introduction
OJ287 is a blazar located at a redshift of z=0.306 ($D_{L}\sim 1600$ Mpc)  (Sitko and Junkkarinen, 1985). This blazar is known to contain a supermassive black hole binary and hence has been a promising target for gravitational wave astronomy (Valtonen et al., 2021; Gao et al., 2024; Chen and Zhang, 2018; Dey et al., 2018; Desai et al., 2008). Optical observations for this object go back to the 19th century (Sillanpaa et al., 1988). This object has been known to emit quasi-periodic outbursts with a period of approximately 12 years. These bursts are emitted due to the enhanced accretion from the secondary black hole on to the primary (Dey et al., 2019; Valtonen et al., 2021, 2023).
In a recent work, Gupta et al. (2023) reported results of the optical flux variability and polarization measurements using a large network of telescopes (Perkins telescope, 1.5 m “Kanata” telescope, St Petersbug University 70-cm AZT-8 telescope and 40-cm LX-200 telescope, 0.6 m telescope Belogradchik Observatory, Steward observatory) spanning eight observing seasons from September 2015 to October 2023. These observations are part of a long-term multiwavelength campaign of OJ 287 (Gupta et al., 2023). Observations were presented for eight segments for the above duration, where each segment roughly corresponds to one observing cycle. The third segment was further subdivided into two sub-segments labelled as 3A and 3B because of some peculiar variability features in the third segment. Most recently,  Valtonen et al. (2024) also reported a huge optical flare in both $I$-band and $R$-band with TESS and the prompt5 telescope in Chile, on Nov 12, 2021 at 2 am UT, which falls during segment 7. Also, TeV gamma-ray emission has also been detected from OJ287 using Veritas between 1 February and 4 February 2017 (UTC) which falls during Segment 2 (VERITAS Collaboration, 2017), although no details on the duration of the TeV flare are available.
In this work, we search for gamma-ray emission during four flaring periods from 2015-2023 using the Fermi Large Area Telescope (LAT) gamma-ray telescope. The LAT is one of the two instruments onboard this detector. Fermi-LAT is sensitive to high-energy gamma rays from a plethora of astrophysical sources (Ajello et al., 2019). It is a pair-conversion telescope with a field of view of aboud 2.4 sr that is sensitive to photons between the energy range of 30 MeV to around 300 GeV and has been operating continuously since 2008 (Atwood et al., 2009). Fermi-LAT has been monitoring OJ287 since the start of the mission. OJ 287 has been included in the 3FGL (Acero et al., 2015) and 4FGL (Abdollahi et al., 2020) catalogs and has been refered to as J0854.8+2006 in the 4FGL (Ajello et al., 2020). Previous gamma-ray variability studies with Fermi-LAT have been carried out in  Prince et al. (2021a). They found that the isotropic gamma-ray luminosity to be between $10^{47}-10^{48}$ ergs/sec. The average and peak photon flux were found to $\sim 2.65$ and $\sim 10$$\times 10^{-8}\rm{ph.cm^{-2}sec^{-1}}$ in the energy range between 0.1-300 GeV (Prince et al., 2021a). A variation of factor of five was found between the lowest and highest flux states from this analysis. We now present the results from our analysis of Fermi-LAT data between 2015-2023.
2 Analysis of Fermi-LAT data
We analyzed the Fermi-LAT data using the EasyFermi software (de Menezes, 2022). The Fermi-LAT data from 2015 - 2023 was divided into eight time segments. For each segment, we searched for OJ 287 in the 4FGL-DR3 catalog, within an energy range of 100 MeV to 300 GeV (unless specified) and a search radius of $7.5^{\circ}$. We used 20-time bins for Light curves for Segments 1, 2, 3A, 5, 7, and 8; 8 bins for Segment 3B; 7 bins for Segment 4; and 10 bins for Segment 6 (as larger number of bins have resulted in null data points in some bins). We then proceeded to fit the Spectral Energy Distribution (SED) using a simple power law model.
$dN/dE=N_{0}\left(E/E_{0}\right)^{-\alpha},$ (1)
where $\alpha$ is the spectral index, $N_{0}$ is the normalization and $E_{0}$ is the pivot energy.
3 Results
We now show the results for the gamma-ray light curves and spectral energy distribution (SED), with and without correction for absorption due to extragalactic background light, and photon flux light curves for OJ 287 during the aforementioned eight segments. A tabular summary of the results is summarized in Table 1, where we show the test statistic (TS) (Mattox et al., 1996; Pasumarti and Desai, 2022; Manna and Desai, 2024), photon flux, energy flux, pivot energy used for normalization, the best-fit spectral parameters and the isotropic gamma-ray luminosity, which has been calculated using Equation 5 of  Prince et al. (2021b).
3.1 Segment 1 (JD: 2457284 - 2457536)
The photon flux light curve (sampled at a cadence of 12 days) for segment 1 and the corresponding SED can be found in Fig. 1 and  2, respectively. The average photon and energy flux are given by $(1.08\pm 0.07)\times 10^{-7}\rm{cm^{-2}s^{-1}}$ and $(4.65\pm 0.27)\times 10^{-5}$ $\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. The light curve shows a variability upto a factor of five between the smallest and largest flux. The SED shows non-zero values up to 20 GeV and upper limits beyond that. The marginalized posteriors for the spectral fit parameters can be found in Fig. 3.
Figure 1: Segment 1 (MJD : 57283.5 - 57535.5 ) - Light Curve
Figure 2: Segment 1 - SED
Figure 3: Segment 1 - SED Best fit parameters: $\log_{10}N_{0}=-9.51\pm 0.03$ ; $\alpha=2.10\pm 0.04$; $E_{0}=636.6$ MeV
3.2 Segment 2 (JD: 2457645 - 2457920)
The photon flux light curve, as well as the SED for segment 2, can be found in Fig. 4 and  5, respectively. The average photon and energy flux are given by $(4.07\pm 0.05)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(3.52\pm 0.09)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. Here, we see a factor of four variation between the smallest and largest flux values. We also find a quasi-periodic variation in the light curves with a period of approximately 12-13 days. The posteriors for the best-fit power law parameters can be found in Fig. 6. The SED shows non-zero flux up to 10 GeV and upper limits beyond that.
Figure 4: Segment 2 (MJD : 57644.5 - 57919.5 ) - Light Curve
Figure 5: Segment 2 - SED
Figure 6: Segment 2 - SED Best fit parameters: $\log_{10}N_{0}=-9.94\pm 0.05$ ; $\alpha=1.92\pm 0.05$ ; $E_{0}=720.59$ MeV
3.3 Segment 3A (JD: 2458016 - 2458142)
The photon flux light curve and SED for segment 3A can be found in Fig. 7 and  8, respectively. The average photon and energy flux are given by $(3.44\pm 0.95)\times\rm{10^{-8}cm^{-2}s^{-1}}$ and $(1.53\pm 0.26)\times\rm{10^{-5}MeV~{}cm^{-2}s^{-1}}$ respectively. Here, we see a nearly constant flux within $1\sigma$ error bars. The SED shows non-zero flux upto 4 GeV and upper limits beyond that. The posteriors for the best-fit power law parameters can be found in Fig. 9.
Figure 7: Segment 3A (MJD : 58015.5 - 58141.5 ) - Light Curve
Figure 8: Segment 3A - SED
Figure 9: Segment 3A - SED Best fit parameters: $\log_{10}N_{0}=-9.89\pm 0.16$ ; $\alpha=2.19\pm 0.19$ ; $E_{0}=846.77$ MeV
3.4 Segment 3B (JD: 2458156 - 2458292)
The photon flux light curve and SED for segment 3B can be found in Fig. 10 and  11, respectively. The average photon and energy flux are given by $(3.88\pm 0.67)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(1.86\pm 0.33)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux until 4 GeV and upper limits beyond that, except for one data point near 50 GeV. Similar to segment 3A, we see an approximately constant flux, except for the first data point, which is close to zero. The posteriors for the best-fit spectral parameters can be found in Fig. 12.
Figure 10: Segment 3B (MJD : $5$ 58291.5 ) - Light Curve
Figure 11: Segment 3B - SED
Figure 12: Segment 3B - SED Best fit parameters: $\log_{10}N_{0}=-10.06\pm 0.12$ ; $\alpha=1.97\pm 0.12$ ; $E_{0}=803.72$ MeV
3.5 Segment 4 (JD: 2458380 - 2458632)
The photon flux light curve and SED for segment 4 can be found in Fig. 13 and  14, respectively. The average photon and energy flux are given by $(4.80\pm 0.78)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.02\pm 0.34)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux until 3 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve shows a signature of quasi-periodic trends, only the period changes from the beginning to the end of the segment. The light curve also shows a factor of five variation between the minimum and maximum flux. The posteriors for the best-fit spectral parameters can be found in Fig. 15.
Figure 13: Segment 4 (MJD : 58379.5 - 58631.5 ) - Light Curve
Figure 14: Segment 4 - SED
Figure 15: Segment 4 - SED Best fit parameters: $\log_{10}N_{0}=-9.90\pm 0.06$ ; $\alpha=2.21\pm 0.10$ ; $E_{0}=545.92$ MeV
3.6 Segment 5 (JD: 2458755 - 2459009)
The photon flux light curve and SED for segment 5 can be found in Fig. 16 and  17, respectively. The average photon and energy flux are given by $(4.52\pm 0.72)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.60\pm 0.45)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux until 20 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve steeply falls near the beginning of the segment (within a factor of 15) and then asymptotes to a nearly constant value for the remainder of the segment. The posteriors for the best-fit spectral parameters can be found in Fig. 18.
Figure 16: Segment 5 (MJD : 58754.5 - 59008.5 ) - Light Curve
Figure 17: Segment 5 - SED
Figure 18: Segment 5 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.05$ ; $\alpha=2.10^{+0.06}_{-0.07}$ ; $E_{0}=921.45$ MeV
3.7 Segment 6 (JD: 2459108 - 2459338)
The photon flux light curve and SED for segment 6 can be found in Fig. 19 and  20, respectively. The average photon and energy flux are given by $(3.72\pm 0.67)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.06\pm 0.33)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 10 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve mainly shows a constant flux, except for the first and third point which depict very small values and then rises towards the end of the segment. The posteriors for the best-fit spectral parameters can be found in Fig. 21.
Figure 19: Segment 6 (MJD : 59107.5 - 59337.5 ) - Light Curve
Figure 20: Segment 6 - SED
Figure 21: Segment 6 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.06$ ; $\alpha=2.04\pm 0.08$ ; $E_{0}=745.69$ MeV
3.8 Segment 7 (JD: 2459520 - 2459750)
The photon flux light curve and SED for segment 7 can be found in Fig. 22 and  23, respectively. The average photon and energy flux are given by $(2.06\pm 0.25)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(1.95\pm 0.27)\times 10^{-5}$ $\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 20 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. The light curve shows a sinusoidal trend near the start of the segment until MJD of 59600 followed by zero flux for the next 50 days. The light curve then shows a constant value until the end of the segment. We also do not find any correlation between the optical flare in $I$ and $R$ band reported in  Valtonen et al. (2024) and the gamma-ray light curve. The posteriors for the best-fit spectral parameters can be found in Fig. 24.
Figure 22: Segment 7 (MJD : 59519.5 - 59749.5 ) - Light Curve. The dashed vertical line corresponds to optical flare detected in $I$ and $R$ band as reported in  Valtonen et al. (2024).
Figure 23: Segment 7 - SED
Figure 24: Segment 7 - SED Best fit parameters: $\log_{10}N_{0}=-10.20^{+0.13}_{-0.12}$ ; $\alpha=1.85^{+0.12}_{-0.11}$ ; $E_{0}=1081.02$ MeV
3.9 Segment 8 (JD: 2459810-2459999)
The photon flux light curve and SED for segment 8 can be found in Fig. 25 and  26, respectively. We note that the optical data was found to be in a low flux state in  Gupta et al. (2023). The average photon and energy flux are given by $(4.61\pm 0.18)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.15\pm 0.14)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. We find that only six data points in the light curve have non-zero values, with the rest showing upper limits or flux values consistent with zero. Similarly the SED only shows values until 2 GeV. The posteriors for the best-fit spectral parameters can be found in Fig. 27.
Figure 25: Segment 8 (MJD : 59809.5 - 59998.5 ) - Light Curve
Figure 26: Segment 8 - SED
Figure 27: Segment 8 - SED Best fit parameters: $\log_{10}N_{0}=-10.07\pm 0.07$ ; $\alpha=1.96\pm 0.08$ ; $E_{0}=705.82$ MeV
3.10 Segments 1 - 8 (JD: 2457284 - 2459999)
We now do a holistic analysis of the data for all the eight segments. The photon flux light curve and SED for segment 8 can be found in Fig. 28 and  29, respectively. The average photon and energy flux are given by $(5.16\pm 0.05)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.71\pm 0.04)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 50 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. We see a falling light curve from the start of segment 1 unil MJD of around 58000. We then see a quasi-sinusoidal trend in the light curves. The posteriors for the best-fit spectral parameters can be found in Fig. 30. The isotropic average gamma-ray luminosity detected during the entire segment is $2.18\times 10^{47}$ ergs/sec.
Figure 28: Segments 1-8 (MJD : 57283.5 - 59998.5 ) - Light Curve
Figure 29: Segments 1-8 - SED
Figure 30: Segments 1-8 - SED Best fit parameters: $\log_{10}N_{0}=-9.87\pm 0.01$ ; $\alpha=2.12\pm 0.02$ ; $E_{0}=715.56$ MeV
Segment TS Flux ×$\times 10^{-8}$ Energy Flux ×$\times 10^{-5}$ $\log(N0)$ $\alpha$ Luminosity ×$\times 10^{47}$ Pivot Energy
(cm-2 s-1) (MeV cm-2 s-1) (erg s-1) $E_{0}$ (MeV)
1 1586.61 $10.79\pm 0.70$ $4.65\pm 0.27$ $-9.51\pm 0.03$ $2.14\pm 0.04$ 3.78 636.6
2 891.89 $4.07\pm 0.05$ $3.52\pm 0.09$ $-9.94\pm 0.05$ $1.92\pm 0.05$ 2.76 720.59
3A 155.29 $3.44\pm 0.95$ $1.53\pm 0.26$ $-9.89\pm 0.16$ $2.19\pm 0.19$ 2.76 846.77
3B 154.6 $3.88\pm 0.67$ $1.86\pm 0.33$ $-10.06\pm 0.12$ $1.97\pm 0.12$ 2.31 803.72
4 263.04 $4.80\pm 0.78$ $2.02\pm 0.34$ $-9.91\pm 0.06$ $2.21\pm 0.10$ 0.98 545.92
5 379.65 $4.52\pm 0.72$ $2.60\pm 0.45$ $-9.96\pm 0.05$ $2.10\pm 0.07$ 3.1 921.45
6 308.48 $3.72\pm 0.67$ $2.06\pm 0.33$ $-9.97\pm 0.06$ $2.03\pm 0.08$ 2.18 745.69
7 320.09 $2.06\pm 0.25$ $1.95\pm 0.27$ $-10.20\pm 0.13$ $1.84\pm 0.12$ 3.96 1081.02
8 173.44 $4.61\pm 0.18$ $2.15\pm 0.14$ $-10.07\pm 0.07$ $1.96\pm 0.08$ 1.79 705.82
all 4995.36 $5.16\pm 0.05$ $2.71\pm 0.04$ $-9.87\pm 0.01$ $2.12\pm 0.02$ 2.18 715.56
Table 1: Summary of test statistics (TS), observed photon and energy flux, pivot energy, $E_{0}$ (cf. Eq 1) along with the best-fit SED parameters for all the eight observing segments. We note that the detection significance is given by the square root of TS (Mattox et al., 1996).
3.11 Comparison with Prince et al
We now compare our results with those in  Prince et al. (2021a). A summary of their observations can be found in Table 4. We find that the spectral index $\alpha$ is in the same ball-park range between 1.85 and 2.24 found in  Prince et al. (2021a). The photon flux in segment 1 is about a factor of two larger than the fluxes detected in Prince et al. (2021a). For all other segments it is comparable. The isotropic luminosity that we detect, which is between $10^{47}-10^{48}$ ergs/sec, is comparable to that detected in  Prince et al. (2021a) and is less than the Eddington luminosity of $10^{50}$ erg/sec. Therefore, our results are mostly consistent with those in  Prince et al. (2021a).
4 Conclusions
In a recent work, Gupta et al. (2023), presented the results of an extensive observing campaign of OJ 287 from 2015-2023 using optical telescopes in three different continents in conjunction with public archival data from the Steward Observatory. These observations were divided into eight segments, where each segment roughly corresponds to an observing season. The third segment was further sub-divided into two sub-segments. We searched for coincident gamma-ray emission from 0.1-300 GeV with the Fermi-LAT telescope during each of the observing segments. For each segment, we show the gamma-ray light curve and SED. We fit the SED to a power law and show a corner plot depicting the best-fit parameters. We also calculate the luminosity for each segment. the plots for the gamma-ray light curves, SED, and marglinalized contours for the SED best-fit parameters for all the eight segments can be found in Figs. 1-30. A tabular summary of our results can be found in Table 1. Our conclusions are as follows:
  • We detect non-zero gamma-ray emission ($>10\sigma$) during all the eight segments, with flux between $(2-10)\times 10^{-8}\rm{cm^{-2}s^{-1}}$
  • The observed energy flux is between $(2-4.6)\times 10^{-5}MeVcm^{-2}s^{-1}$.
  • Ths isotropic luminosity is between $(0.98-3.96)\times 10^{47}$ ergs/sec.
  • The power-law index varies between 1.8 - 2.2.
  • The maximum flux is seen near the start of first segment. The flux decreases unil MJD of 58000.
  • We do not find any correlated gamma-ray flare during Segment 7 in coincidence with an optical flare in $I$ and $R$ band recently reported in  Valtonen et al. (2024).
  • We do not detect any emission beyond 20-30 GeV.
Acknowledgments
We are grateful to Raniere Menezes for help with the EasyFermi package.

### v2

@@
 We present results for eight segments in the aforementioned period, with each segment corresponding to an observing season.
-We find non-zero gamma-ray emission for all the eight segments.
+We report non-zero gamma-ray flux (with $>10\sigma$ significance) for all the eight segments.
 The photon and energy flux observed during this period is $\sim 5.2\times 10^{-8}\rm{ph~{}cm^{-2}~{}s^{-1}}$ and $2.75\times 10^{-5}\rm{MeV~{}cm^{-2}~{}s^{-1}}$, respectively, while the SED can be fitted with a power law with a slope of $2.12\pm 0.02$.
 The observed luminosity is between $10^{45}-10^{46}$ ergs/sec.
+Similar to previous works we find a hard cutoff in the gamma-ray spectrum at around 20 GeV.
+We do not observe any correlations between the peaks in the gamma ray light curves and corresponding optical light curves in $I$ and $R$ bands, as well as the X-ray light curves from SWIFT between 0.3-10 keV.
 1 Introduction
-OJ287 is a blazar located at a redshift of z=0.306 ($D_{L}\sim 1600$ Mpc)  (Sitko and Junkkarinen, 1985).
-This blazar is known to contain a supermassive black hole binary and hence has been a promising target for gravitational wave astronomy (Valtonen et al., 2021; Gao et al., 2024; Chen and Zhang, 2018; Dey et al., 2018; Desai et al., 2008).
-Optical observations for this object go back to the 19th century (Sillanpaa et al., 1988).
+OJ287 is a BL object located at a redshift of z=0.306 ($D_{L}\sim 1600$ Mpc)  (Sitko and Junkkarinen, 1985).
+This object is known to contain a supermassive black hole binary and hence has been a promising target for studies in gravitational wave astronomy (Valtonen et al., 2021; Gao et al., 2024; Chen and Zhang, 2018; Dey et al., 2018; Desai et al., 2008).
+Optical observations for this object go back to the 19th century (Sillanpaa et al., 1988; Hudec et al., 2013).
 This object has been known to emit quasi-periodic outbursts with a period of approximately 12 years.
-These bursts are emitted due to the enhanced accretion from the secondary black hole on to the primary (Dey et al., 2019; Valtonen et al., 2021, 2023).
+These optical bursts are emitted due to the enhanced accretion from the secondary black hole onto the primary (Dey et al., 2019; Valtonen et al., 2021, 2023).
 In a recent work, Gupta et al.
-(2023) reported results of the optical flux variability and polarization measurements using a large network of telescopes (Perkins telescope, 1.5 m “Kanata” telescope, St Petersbug University 70-cm AZT-8 telescope and 40-cm LX-200 telescope, 0.6 m telescope Belogradchik Observatory, Steward observatory) spanning eight observing seasons from September 2015 to October 2023.
+(2023) reported results related to optical flux variability and polarization measurements using a large network of telescopes (Perkins telescope, 1.5 m “Kanata” telescope, St Petersburg University 70-cm AZT-8 telescope and 40-cm LX-200 telescope, 0.6 m telescope at the Belogradchik Observatory, Steward Observatory telescope) spanning eight observing seasons from September 2015 to October 2023.
 These observations are part of a long-term multiwavelength campaign of OJ 287 (Gupta et al., 2023).
@@
 Most recently,  Valtonen et al.
-(2024) also reported a huge optical flare in both $I$-band and $R$-band with TESS and the prompt5 telescope in Chile, on Nov 12, 2021 at 2 am UT, which falls during segment 7.
-Also, TeV gamma-ray emission has also been detected from OJ287 using Veritas between 1 February and 4 February 2017 (UTC) which falls during Segment 2 (VERITAS Collaboration, 2017), although no details on the duration of the TeV flare are available.
-In this work, we search for gamma-ray emission during four flaring periods from 2015-2023 using the Fermi Large Area Telescope (LAT) gamma-ray telescope.
+(2024) also reported a huge optical flare in November 2021 in both $I$-band and $R$-band with TESS and the prompt5 telescope in Chile as part of the Krakow quasar monitoring campaign, which falls during segment 7.
+Also, TeV gamma-ray emission has also been detected from OJ287 using VERITAS between 1 February and 4 February 2017 (UTC), which falls during Segment 2 (VERITAS Collaboration, 2017), although no details on the duration of the TeV flare are available.
+Finally this object has also been monitored in X-rays part of the MOMO project (Komossa et al., 2023a).
+In this work, we search for gamma-ray emission during the eight aforementioned segments from 2015-2023 using the Fermi Large Area Telescope (LAT) gamma-ray telescope.
 The LAT is one of the two instruments onboard this detector.
 Fermi-LAT is sensitive to high-energy gamma rays from a plethora of astrophysical sources (Ajello et al., 2019).
-It is a pair-conversion telescope with a field of view of aboud 2.4 sr that is sensitive to photons between the energy range of 30 MeV to around 300 GeV and has been operating continuously since 2008 (Atwood et al., 2009).
+It is a pair-conversion telescope with a field of view of about 2.4 sr that is sensitive to photons between the energy range of 30 MeV to around 300 GeV and has been operating continuously since 2008 (Atwood et al., 2009).
 Fermi-LAT has been monitoring OJ287 since the start of the mission.
-OJ 287 has been included in the 3FGL (Acero et al., 2015) and 4FGL (Abdollahi et al., 2020) catalogs and has been refered to as J0854.8+2006 in the 4FGL (Ajello et al., 2020).
-Previous gamma-ray variability studies with Fermi-LAT have been carried out in  Prince et al.
-(2021a).
-They found that the isotropic gamma-ray luminosity to be between $10^{47}-10^{48}$ ergs/sec.
-The average and peak photon flux were found to $\sim 2.65$ and $\sim 10$$\times 10^{-8}\rm{ph.cm^{-2}sec^{-1}}$ in the energy range between 0.1-300 GeV (Prince et al., 2021a).
-A variation of factor of five was found between the lowest and highest flux states from this analysis.
+OJ 287 has been included in the 3FGL (Acero et al., 2015) and 4FGL (Abdollahi et al., 2020) catalogs and has been referred to as J0854.8+2006 in the 4FGL (Ajello et al., 2020).
+Previous gamma-ray variability studies with Fermi-LAT have been reported in a number of works (Abdo et al., 2009; Neronov and Vovk, 2011; Kushwaha et al., 2013; Prince et al., 2021a; Komossa et al., 2022).
+Prince et al.
+(2021a) found that the isotropic gamma-ray luminosity to be between $10^{47}-10^{48}$ ergs/sec.
+The average and peak photon flux were found to be $\sim 2.65$ and $\sim 10$$\times 10^{-8}\rm{ph.cm^{-2}sec^{-1}}$ in the energy range between 0.1-300 GeV (Prince et al., 2021a).
+This analysis found a variation of a factor of five between the lowest and highest flux states.
 We now present the results from our analysis of Fermi-LAT data between 2015-2023.
+This manuscript is structured as follows.
+Our analysis of the Fermi-LAT data is discussed in Sect 2.
+Comparison with other multi-wavelength flares for OJ 287 is discussed in Sect.
+4.2.
+We conclude in Sect.
+5.
 2 Analysis of Fermi-LAT data
-We analyzed the Fermi-LAT data using the EasyFermi software (de Menezes, 2022).
+We analyzed the Fermi-LAT data using the EasyFermi software (de Menezes, 2022), which is based on both FermiTools and Fermipy packages (Wood et al., 2017).
 The Fermi-LAT data from 2015 - 2023 was divided into eight time segments.
@@
 We used 20-time bins for light curves for Segments 1, 2, 3A, 5, 7, and 8; 8 bins for Segment 3B; 7 bins for Segment 4; and 10 bins for Segment 6 (as larger number of bins have resulted in null data points in some bins).
-We then proceeded to fit the Spectral Energy Distribution (SED) using a simple power law model.
+We then proceeded to fit the Spectral Energy Distribution (SED) using a simple power law model and also a log-parabola model.
+The power-law spectrum is defined as follows:
 $dN/dE=N_{0}\left(E/E_{0}\right)^{-\alpha},$ (1)
-where $\alpha$ is the spectral index, $N_{0}$ is the normalization and $E_{0}$ is the pivot energy.
+where $\alpha$ is the spectral index, $N_{0}$ is the normalization and $E_{0}$ is the pivot energy. Similarly, the log-parabola spectrum is defined according to:
+$dN/dE=N_{0}\left(\frac{E}{E_{0}}\right)^{-\alpha-\beta\log{\left(\frac{E}{E_{0}}\right)}}$ (2)
 3 Results
-We now show the results for the gamma-ray light curves and spectral energy distribution (SED), with and without correction for absorption due to extragalactic background light, and photon flux light curves for OJ 287 during the aforementioned eight segments.
-A tabular summary of the results is summarized in Table 1, where we show the test statistic (TS) (Mattox et al., 1996; Pasumarti and Desai, 2022; Manna and Desai, 2024), photon flux, energy flux, pivot energy used for normalization, the best-fit spectral parameters and the isotropic gamma-ray luminosity, which has been calculated using Equation 5 of  Prince et al.
+We now show the results for the gamma-ray light curves and SED, with and without correction for absorption due to extragalactic background light, and photon flux light curves for OJ 287 during the aforementioned eight segments.
+A tabular summary of the results is summarized in Table 1, where we show the test statistic (TS) (Mattox et al., 1996; Pasumarti and Desai, 2022; Manna and Desai, 2024), photon flux, energy flux, and the isotropic gamma-ray luminosity, which has been calculated using Equation 5 of  Prince et al.
 (2021b).
+The best-fit spectral parameters for both the power-law and log-parabola model can be found in Table 2.
+We also confirmed that that the value of “Fit Quality” reported by EasyFermi is equal to three for all the analyzed segments, indicating the fit is excellent and there is no problem in the convergence of the log likelihood de Menezes (2022).
 3.1 Segment 1 (JD: 2457284 - 2457536)
@@
 3.
-Figure 1: Segment 1 (MJD : 57283.5 - 57535.5 ) - Light Curve
-Figure 2: Segment 1 - SED
-Figure 3: Segment 1 - SED Best fit parameters: $\log_{10}N_{0}=-9.51\pm 0.03$ ; $\alpha=2.10\pm 0.04$; $E_{0}=636.6$ MeV
+Figure 1: Segment 1 (MJD : 57283.5 - 57535.5 ) - Light Curve; binwidth = 12 days.
+Figure 2: Segment 1 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 3: Segment 1 - SED Best fit parameters: $\log_{10}N_{0}=-9.51\pm 0.03$ ; $\alpha=2.10\pm 0.04$; $E_{0}=636.6$ MeV.
 3.2 Segment 2 (JD: 2457645 - 2457920)
@@
 Here, we see a factor of four variation between the smallest and largest flux values.
-We also find a quasi-periodic variation in the light curves with a period of approximately 12-13 days.
 The posteriors for the best-fit power law parameters can be found in Fig.
@@
 The SED shows non-zero flux up to 10 GeV and upper limits beyond that.
-Figure 4: Segment 2 (MJD : 57644.5 - 57919.5 ) - Light Curve
-Figure 5: Segment 2 - SED
-Figure 6: Segment 2 - SED Best fit parameters: $\log_{10}N_{0}=-9.94\pm 0.05$ ; $\alpha=1.92\pm 0.05$ ; $E_{0}=720.59$ MeV
+Figure 4: Segment 2 (MJD : 57644.5 - 57919.5 ) - Light Curve; binwidth = 13.75 days
+Figure 5: Segment 2 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 6: Segment 2 - SED Best fit parameters: $\log_{10}N_{0}=-9.94\pm 0.05$ ; $\alpha=1.92\pm 0.05$ ; $E_{0}=720.59$ MeV.
 3.3 Segment 3A (JD: 2458016 - 2458142)
@@
 9.
-Figure 7: Segment 3A (MJD : 58015.5 - 58141.5 ) - Light Curve
-Figure 8: Segment 3A - SED
-Figure 9: Segment 3A - SED Best fit parameters: $\log_{10}N_{0}=-9.89\pm 0.16$ ; $\alpha=2.19\pm 0.19$ ; $E_{0}=846.77$ MeV
+Figure 7: Segment 3A (MJD : 58015.5 - 58141.5 ) - Light Curve; binwidth = 6.3 days
+Figure 8: Segment 3A - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 9: Segment 3A - SED Best fit parameters: $\log_{10}N_{0}=-9.89\pm 0.16$ ; $\alpha=2.19\pm 0.19$ ; $E_{0}=846.77$ MeV.
 3.4 Segment 3B (JD: 2458156 - 2458292)
@@
 From the SED, we find non-zero flux until 4 GeV and upper limits beyond that, except for one data point near 50 GeV.
-Similar to segment 3A, we see an approximately constant flux, except for the first data point, which is close to zero.
+Similar to segment 3A, we see an approximately constant flux, except for the first data point.
 The posteriors for the best-fit spectral parameters can be found in Fig.
 12.
-Figure 10: Segment 3B (MJD : $5$ 58291.5 ) - Light Curve
-Figure 11: Segment 3B - SED
-Figure 12: Segment 3B - SED Best fit parameters: $\log_{10}N_{0}=-10.06\pm 0.12$ ; $\alpha=1.97\pm 0.12$ ; $E_{0}=803.72$ MeV
+Figure 10: Segment 3B (MJD : $5$ 58291.5 ) - Light Curve ; binwidth = 17 days
+Figure 11: Segment 3B - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 12: Segment 3B - SED Best fit parameters: $\log_{10}N_{0}=-10.06\pm 0.12$ ; $\alpha=1.97\pm 0.12$ ; $E_{0}=803.72$ MeV.
 3.5 Segment 4 (JD: 2458380 - 2458632)
@@
 Similar to segment 3A, we see an approximately constant flux, except for the first data point.
-The light curve shows a signature of quasi-periodic trends, only the period changes from the beginning to the end of the segment.
 The light curve also shows a factor of five variation between the minimum and maximum flux.
@@
 15.
-Figure 13: Segment 4 (MJD : 58379.5 - 58631.5 ) - Light Curve
-Figure 14: Segment 4 - SED
-Figure 15: Segment 4 - SED Best fit parameters: $\log_{10}N_{0}=-9.90\pm 0.06$ ; $\alpha=2.21\pm 0.10$ ; $E_{0}=545.92$ MeV
+Figure 13: Segment 4 (MJD : 58379.5 - 58631.5 ) - Light Curve ; binwidth = 36 days
+Figure 14: Segment 4 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 15: Segment 4 - SED Best fit parameters: $\log_{10}N_{0}=-9.90\pm 0.06$ ; $\alpha=2.21\pm 0.10$ ; $E_{0}=545.92$ MeV.
 3.6 Segment 5 (JD: 2458755 - 2459009)
@@
 18.
-Figure 16: Segment 5 (MJD : 58754.5 - 59008.5 ) - Light Curve
-Figure 17: Segment 5 - SED
-Figure 18: Segment 5 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.05$ ; $\alpha=2.10^{+0.06}_{-0.07}$ ; $E_{0}=921.45$ MeV
+Figure 16: Segment 5 (MJD : 58754.5 - 59008.5 ) - Light Curve ; binwidth = 12.7 days
+Figure 17: Segment 5 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 18: Segment 5 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.05$ ; $\alpha=2.10^{+0.06}_{-0.07}$ ; $E_{0}=921.45$ MeV.
 3.7 Segment 6 (JD: 2459108 - 2459338)
@@
 Similar to segment 3A, we see an approximately constant flux, except for the first data point.
-The light curve mainly shows a constant flux, except for the first and third point which depict very small values and then rises towards the end of the segment.
+The light curve mainly shows a constant flux, except for the first and third points, which depict very small values and then rise towards the end of the segment.
 The posteriors for the best-fit spectral parameters can be found in Fig.
 21.
-Figure 19: Segment 6 (MJD : 59107.5 - 59337.5 ) - Light Curve
-Figure 20: Segment 6 - SED
-Figure 21: Segment 6 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.06$ ; $\alpha=2.04\pm 0.08$ ; $E_{0}=745.69$ MeV
+Figure 19: Segment 6 (MJD : 59107.5 - 59337.5 ) - Light Curve ; binwidth = 23 days
+Figure 20: Segment 6 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 21: Segment 6 - SED Best fit parameters: $\log_{10}N_{0}=-9.96\pm 0.06$ ; $\alpha=2.04\pm 0.08$ ; $E_{0}=745.69$ MeV.
 3.8 Segment 7 (JD: 2459520 - 2459750)
@@
 Similar to segment 3A, we see an approximately constant flux, except for the first data point.
-The light curve shows a sinusoidal trend near the start of the segment until MJD of 59600 followed by zero flux for the next 50 days.
+The light curve shows a sinusoidal trend near the start of the segment until MJD of 59600 followed by upper limits for the next 50 days.
 The light curve then shows a constant value until the end of the segment.
@@
 24.
-Figure 22: Segment 7 (MJD : 59519.5 - 59749.5 ) - Light Curve. The dashed vertical line corresponds to optical flare detected in $I$ and $R$ band as reported in  Valtonen et al. (2024).
-Figure 23: Segment 7 - SED
-Figure 24: Segment 7 - SED Best fit parameters: $\log_{10}N_{0}=-10.20^{+0.13}_{-0.12}$ ; $\alpha=1.85^{+0.12}_{-0.11}$ ; $E_{0}=1081.02$ MeV
+Figure 22: Segment 7 (MJD : 59519.5 - 59749.5 ) - Light Curve ; binwidth = 11.5 days.
+The dashed vertical line corresponds to the start of the optical flare detected in $I$ and $R$ band as reported in  Valtonen et al.
+(2024).
+Figure 23: Segment 7 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 24: Segment 7 - SED Best fit parameters: $\log_{10}N_{0}=-10.20^{+0.13}_{-0.12}$ ; $\alpha=1.85^{+0.12}_{-0.11}$ ; $E_{0}=1081.02$ MeV.
 3.9 Segment 8 (JD: 2459810-2459999)
@@
 The average photon and energy flux are given by $(4.61\pm 0.18)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.15\pm 0.14)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively.
-We find that only six data points in the light curve have non-zero values, with the rest showing upper limits or flux values consistent with zero.
-Similarly the SED only shows values until 2 GeV.
+We find that only six data points in the light curve have non-zero values, with the rest showing upper limits.
+Similarly, the SED only shows values until 2 GeV.
 The posteriors for the best-fit spectral parameters can be found in Fig.
 27.
-Figure 25: Segment 8 (MJD : 59809.5 - 59998.5 ) - Light Curve
-Figure 26: Segment 8 - SED
-Figure 27: Segment 8 - SED Best fit parameters: $\log_{10}N_{0}=-10.07\pm 0.07$ ; $\alpha=1.96\pm 0.08$ ; $E_{0}=705.82$ MeV
+Figure 25: Segment 8 (MJD : 59809.5 - 59998.5 ) - Light Curve ; binwidth = 9.45 days.
+Figure 26: Segment 8 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 27: Segment 8 - SED Best fit parameters: $\log_{10}N_{0}=-10.07\pm 0.07$ ; $\alpha=1.96\pm 0.08$ ; $E_{0}=705.82$ MeV.
 3.10 Segments 1 - 8 (JD: 2457284 - 2459999)
-We now do a holistic analysis of the data for all the eight segments.
+We now do a holistic analysis of the entire data for all eight segments.
 The photon flux light curve and SED for segment 8 can be found in Fig.
@@
 The average photon and energy flux are given by $(5.16\pm 0.05)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.71\pm 0.04)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively.
-From the SED, we find non-zero flux up to approximately 50 GeV and upper limits beyond that.
+From the SED, we find non-zero flux up to approximately 20 GeV and upper limits beyond that.
 Similar to segment 3A, we see an approximately constant flux, except for the first data point.
-We see a falling light curve from the start of segment 1 unil MJD of around 58000.
-We then see a quasi-sinusoidal trend in the light curves.
+We see a falling light curve from the start of segment 1 until MJD of around 58000.
+We also confirmed that we do not find any signal beyond 20 GeV even aftercoarser binning.
 The posteriors for the best-fit spectral parameters can be found in Fig.
@@
 The isotropic average gamma-ray luminosity detected during the entire segment is $2.18\times 10^{47}$ ergs/sec.
-Figure 28: Segments 1-8 (MJD : 57283.5 - 59998.5 ) - Light Curve
-Figure 29: Segments 1-8 - SED
-Figure 30: Segments 1-8 - SED Best fit parameters: $\log_{10}N_{0}=-9.87\pm 0.01$ ; $\alpha=2.12\pm 0.02$ ; $E_{0}=715.56$ MeV
-Segment TS Flux ×$\times 10^{-8}$ Energy Flux ×$\times 10^{-5}$ $\log(N0)$ $\alpha$ Luminosity ×$\times 10^{47}$ Pivot Energy
-(cm-2 s-1) (MeV cm-2 s-1) (erg s-1) $E_{0}$ (MeV)
-1 1586.61 $10.79\pm 0.70$ $4.65\pm 0.27$ $-9.51\pm 0.03$ $2.14\pm 0.04$ 3.78 636.6
-2 891.89 $4.07\pm 0.05$ $3.52\pm 0.09$ $-9.94\pm 0.05$ $1.92\pm 0.05$ 2.76 720.59
-3A 155.29 $3.44\pm 0.95$ $1.53\pm 0.26$ $-9.89\pm 0.16$ $2.19\pm 0.19$ 2.76 846.77
-3B 154.6 $3.88\pm 0.67$ $1.86\pm 0.33$ $-10.06\pm 0.12$ $1.97\pm 0.12$ 2.31 803.72
-4 263.04 $4.80\pm 0.78$ $2.02\pm 0.34$ $-9.91\pm 0.06$ $2.21\pm 0.10$ 0.98 545.92
-5 379.65 $4.52\pm 0.72$ $2.60\pm 0.45$ $-9.96\pm 0.05$ $2.10\pm 0.07$ 3.1 921.45
-6 308.48 $3.72\pm 0.67$ $2.06\pm 0.33$ $-9.97\pm 0.06$ $2.03\pm 0.08$ 2.18 745.69
-7 320.09 $2.06\pm 0.25$ $1.95\pm 0.27$ $-10.20\pm 0.13$ $1.84\pm 0.12$ 3.96 1081.02
-8 173.44 $4.61\pm 0.18$ $2.15\pm 0.14$ $-10.07\pm 0.07$ $1.96\pm 0.08$ 1.79 705.82
-all 4995.36 $5.16\pm 0.05$ $2.71\pm 0.04$ $-9.87\pm 0.01$ $2.12\pm 0.02$ 2.18 715.56
-Table 1: Summary of test statistics (TS), observed photon and energy flux, pivot energy, $E_{0}$ (cf.
-Eq 1) along with the best-fit SED parameters for all the eight observing segments.
+We also did a search for periodicity using the data for all eight segments using the generalized Lomb-Scargle periodogram (VanderPlas, 2018).
+For this purpose, we use the same implementation as our most recent work on the analysis of Super-Kamiokande solar neutrino data (Pasumarti and Desai, 2024).
+We searched using a frequency grid with a resolution of 0.028 $yr^{-1}$ and ranging from 0 to 2.69 $yr^{-1}$.
+The largest LS power is at 2.15 $yr^{-1}$.
+However, the $p$-value using all the metrics specified in  VanderPlas (2018) is greater than 0.1.
+Therefore, we conclude that there is no discernible periodicity in the OJ 287 gamma-ray dataset during the above period.
+Figure 28: Segments 1-8 (MJD : 57283.5 - 59998.5) - Light Curve ; binwidth = 135.75 days.
+Figure 29: Segments 1-8 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
+Figure 30: Segments 1-8 - SED Best fit parameters: $\log_{10}N_{0}=-9.87\pm 0.01$ ; $\alpha=2.12\pm 0.02$ ; $E_{0}=715.56$ MeV.
+Segment TS Flux ×$\times 10^{-8}$ Energy Flux ×$\times 10^{-5}$ Luminosity
+(cm-2 s-1) (MeV cm-2 s-1) ×$\times 10^{47}$ (erg s-1)
+1 1586.61 $10.79\pm 0.70$ $4.65\pm 0.27$ 3.78
+2 891.89 $4.07\pm 0.05$ $3.52\pm 0.09$ 2.76
+3A 155.29 $3.44\pm 0.95$ $1.53\pm 0.26$ 2.76
+3B 154.6 $3.88\pm 0.67$ $1.86\pm 0.33$ 2.31
+4 263.04 $4.80\pm 0.78$ $2.02\pm 0.34$ 0.98
+5 379.65 $4.52\pm 0.72$ $2.60\pm 0.45$ 3.1
+6 308.48 $3.72\pm 0.67$ $2.06\pm 0.33$ 2.18
+7 320.09 $2.06\pm 0.25$ $1.95\pm 0.27$ 3.96
+8 173.44 $4.61\pm 0.18$ $2.15\pm 0.14$ 1.79
+all 4995.36 $5.16\pm 0.05$ $2.71\pm 0.04$ 2.18
+Table 1: Summary of test statistics (TS), observed photon, energy flux, and the isotropic gamma-ray luminosity.
 We note that the detection significance is given by the square root of TS (Mattox et al., 1996).
-3.11 Comparison with Prince et al
-We now compare our results with those in  Prince et al.
+Simple PL fit Log Parabola fit
+Segment Pivot $E_{0}$ (MeV) $\log_{10}$0 $\alpha$ $\log_{10}$0 $\alpha$ $\beta$ $E_{b}$ (MeV)
+1 $-9.51\pm 0.03$ $2.14\pm 0.04$ 636.6 $-10.65\pm 0.04$ $2.15\pm 0.04$ $0.10\pm 0.02$ 710.99
+2 $-9.94\pm 0.05$ $1.92\pm 0.05$ 720.59 $-10.97\pm 0.01$ $1.80\pm 0.01$ $0.07\pm 0.00$ 710.99
+3A $-9.89\pm 0.16$ $2.19\pm 0.19$ 846.77 $-11.10\pm 0.13$ $2.11\pm 0.16$ $0.14\pm 0.08$ 710.99
+3B $-10.06\pm 0.12$ $1.97\pm 0.12$ 803.72 $-11.16\pm 0.10$ $2.15\pm 0.09$ $0.03\pm 0.04$ 710.99
+4 $-9.91\pm 0.06$ $2.21\pm 0.10$ 545.92 $-11.15\pm 0.12$ $2.27\pm 0.09$ $0.00\pm 0.04$ 710.99
+5 $-9.96\pm 0.05$ $2.10\pm 0.07$ 921.45 $-11.11\pm 0.09$ $2.13\pm 0.08$ $0.00\pm 0.03$ 710.99
+6 $-9.97\pm 0.06$ $2.03\pm 0.08$ 745.69 $-11.11\pm 0.11$ $2.04\pm 0.10$ $0.05\pm 0.05$ 710.99
+7 $-10.20\pm 0.13$ $1.84\pm 0.12$ 1081.02 $-11.19\pm 0.09$ $1.69\pm 0.07$ $0.12\pm 0.03$ 710.99
+8 $-10.07\pm 0.07$ $1.96\pm 0.08$ 705.82 $-11.26\pm 0.03$ $2.33\pm 0.02$ $-0.07\pm 0.01$ 710.99
+all $-9.87\pm 0.01$ $2.12\pm 0.02$ 715.56 $-11.02\pm 0.01$ $2.11\pm 0.01$ $0.03\pm 0.00$ 710.99
+Table 2: Results of spectral fits using the power-law (cf. Eq. 1) as well as log-parabola model (cf. Eq. 2) for all the eight segments
+4 Comparison with other monitoring programs for OJ 287
+4.1 Comparison with Prince et al.
+We now compare our results with the gamma-ray observations for OJ 287 reported in  Prince et al.
 (2021a).
@@
 (2021a).
-For all other segments it is comparable.
+For all other segments, it is comparable.
 The isotropic luminosity that we detect, which is between $10^{47}-10^{48}$ ergs/sec, is comparable to that detected in  Prince et al.
@@
 (2021a).
-4 Conclusions
+Figure 31: Comparison of the optical $R$-band observations reported in  Valtonen et al.
+(2024) alongside the flux observed by Fermi-LAT during Segment 7.
+The gamma-ray and optical light curves have been normalized by their respective modes.
+4.2 Comparison with $R$-band optical flare in 2021-2022
+Most recently,  Valtonen et al.
+(2024) reported optical flares in the $R$-band between October 21st and November 2021 as part of the Krakow quasar monitoring campaign consisting of several telescopes located all over the globe.
+These flares are caused by the impact of the secondary black hole onto the accretion disk of the primary black hole.
+This flare occurred during segment 7.
+To compare the optical data with the Fermi-LAT data, we plot both in the same figure after normalizing each data by the mode.
+This plot can be found in Fig.
+31.
+As we can see, although the gamma-ray light curve is sampled coarsely compared to the optical light curve, there is no direct one-to-one correlation between the rise and fall time of the gamma-ray and optical light curves during this segment.
+The fall in the optical light curve coincides with the rising gamma-ray flux at the beginning of the segment, where the peaks in the two light curves are separated by around 200 days.
+4.3 Comparison with SWIFT XRT light curves from 2015-2021
+OJ 287 has also been monitored by the SWIFT satellite in optical/UV bands using six filters as well as X-rays from 0.3-10 keV as part of the MOMO (Multiwavelength Observations and Modelling of OJ 287) project (Komossa et al., 2023a) from September 2015 till September 2021.
+Additional SWIFT observations starting from 2015 have also been reported (Komossa et al., 2020, 2021).
+These observations have been collated in  Komossa et al.
+(2022).
+Theoretical implications of the observations from the MOMO project can be found in  Komossa et al.
+(2023b, c).
+We overlay these SWIFT X-ray observations along with the Fermi-LAT observations in Fig.
+32.
+Once again we do not see any correlation between the peaks in the SWIFT XRT light curves and gamma-ray light curves.
+Figure 32: Comparison of the SWIFT XRT observations from 0.3-10 keV obtained from Komossa et al.
+(2022), alongside the flux observed by Fermi-LAT.
+The gamma-ray and X-ray observations are normalized by their respective modes.
+5 Discussion and Conclusions
 In a recent work, Gupta et al.
-(2023), presented the results of an extensive observing campaign of OJ 287 from 2015-2023 using optical telescopes in three different continents in conjunction with public archival data from the Steward Observatory.
-These observations were divided into eight segments, where each segment roughly corresponds to an observing season.
-The third segment was further sub-divided into two sub-segments.
-We searched for coincident gamma-ray emission from 0.1-300 GeV with the Fermi-LAT telescope during each of the observing segments.
+(2023) presented the results of an extensive observing campaign of OJ 287 from 2015-2023 using optical telescopes in three different continents in conjunction with public archival data from the Steward Observatory.
+These observations were divided into eight segments, each roughly corresponding to an observing season.
+The third segment was further subdivided into two sub-segments.
+We searched for coincident gamma-ray emission from 0.1-300 GeV with the Fermi-LAT telescope during each observing segment.
 For each segment, we show the gamma-ray light curve and SED.
@@
 We also calculate the luminosity for each segment.
-the plots for the gamma-ray light curves, SED, and marglinalized contours for the SED best-fit parameters for all the eight segments can be found in Figs.
+the plots for the gamma-ray light curves, SED, and marginalized contours for the SED best-fit parameters for all the eight segments can be found in Figs.
 1-30.
-A tabular summary of our results can be found in Table 1.
+Tables  1 and  2 provides a tabular summary of our results.
 Our conclusions are as follows:
   • We detect non-zero gamma-ray emission ($>10\sigma$) during all the eight segments, with flux between $(2-10)\times 10^{-8}\rm{cm^{-2}s^{-1}}$
-  • The observed energy flux is between $(2-4.6)\times 10^{-5}MeVcm^{-2}s^{-1}$.
+  • The observed energy flux is between $(2-4.6)\times 10^{-5}~{}\rm{MeVcm^{-2}s^{-1}}$.
   • Ths isotropic luminosity is between $(0.98-3.96)\times 10^{47}$ ergs/sec.
   • The power-law index varies between 1.8 - 2.2.
-  • The maximum flux is seen near the start of first segment. The flux decreases unil MJD of 58000.
-  • We do not find any correlated gamma-ray flare during Segment 7 in coincidence with an optical flare in $I$ and $R$ band recently reported in  Valtonen et al. (2024).
-  • We do not detect any emission beyond 20-30 GeV.
+  • The maximum flux is seen near the start of the first segment. The flux decreases until MJD of 58000.
+  • We do not find any correlated gamma-ray flare during Segment 7 in coincidence with an optical flare in $I$ and $R$ band recently reported in  Valtonen et al.
+(2024).
+A comparison of the gamma-ray and optical light curves can be found in Fig.
+31.
+  • We also do not find any correlation between the peaks in the gamma-ray flux and X-ray flux observed by the SWIFT satellite between 0.3-10 keV (cf. Fig 32).
+  • We do not detect any emission beyond 20-30 GeV, similar to previously reported gamma-ray observation from OJ287 (Prince et al., 2021a; Kushwaha et al., 2013).
+The gamma-ray spectrum in OJ287 has been previously explained due to a combination of Synchrotron self-Compton and external Compton processes from the scattering of soft photons that are outside the jet (Kushwaha et al., 2013).
+The cut-off in the spectrum can be explained with a power law distribution of electrons with a maximum energy of approximately $10^{4}$ times the electron rest mass Kushwaha et al.
+(2013).
 Acknowledgments
-We are grateful to Raniere Menezes for help with the EasyFermi package.
+We are grateful to Raniere Menezes for his help with the EasyFermi package.
+We are grateful to Stefanie Komossa for providing us the SWIFT X-ray data from  Komossa et al.
+(2022) and Mauri Valotnen for $R$-band data reported in  Valtonen et al.
+(2024).
+We also acknowledge the anonymous referee for constructive feedback on our manuscript.